# Logical forms complement probability in understanding language model (and human) performance

### v1

@@ Abstract
With the increasing interest in using large language models (LLMs) for planning in natural language, understanding their behaviors becomes an important research question. This work conducts a systematic investigation of LLMs’ ability to perform logical reasoning in natural language. We introduce a controlled dataset of hypothetical and disjunctive syllogisms in propositional and modal logic and use it as the testbed for understanding LLM performance. Our results lead to novel insights in predicting LLM behaviors: in addition to the probability of input (Gonen et al., 2023; McCoy et al., 2024), logical forms should be considered as orthogonal factors. In addition, we show similarities and differences between the logical reasoning performances of humans and LLMs by comparing LLM and human behavioral results.
@@ 1 Introduction
In this work, we investigate the logical reasoning capabilities of LLMs by assessing their performance on different logical forms. We curate a dataset of hypothetical syllogisms and disjunctive syllogisms in propositional and modal logic, where some logical forms are semantically equivalent to others. These forms are then interpreted with meaningful real-world statements that mirror the reasoning in daily communication, or phrases that consist of nonsensical words (Figure 2).
We conduct a series of controlled experiments to analyze the performance of a set of LLMs on these questions. Although our findings generally align with those by Gonen et al. (2023) and McCoy et al. (2024), who suggest that LLMs excel on examples with high probability, our results indicate that logical form, including but not limited to modalities and argument forms, is a crucial complementary factor in predicting the performance of LLMs (Figure 1). Additionally, with meaningful real-world interpretations, we find that:
@@
This paper is structured as follows. After reviewing related literature (§ 2), we describe the dataset synthesis process (§ 3). We report the LLM reasoning results on the synthesized dataset (§ 4) and ground the LLM reasoning performance to human reasoning performance (§ 5). We conclude by discussing the implications of our results and the limitations of this work (§ 6).
@@ 2 Related Work
Logical reasoning benchmarks. Prior LLM logical reasoning benchmarks (Liu et al., 2020; Han et al., 2022, inter alia) focus on complex, multi-hop reasoning problems with manually annotated problems, making cross-problem comparisons challenging. Recent work has introduced benchmarks with synthesized natural-language questions using predefined logical formulas and substitution rules (Saparov and He, 2022; Saparov et al., 2023; Parmar et al., 2024; Wan et al., 2024, inter alia). Compared to them, our work uniquely incorporates modal logic, which has been largely unexplored in existing benchmarks—while Holliday and Mandelkern (2024) present a case study, our approach offers two key advances: controlled knowledge bias in logic interpretations (§ 3.3) and a more rigorous statistical evaluation framework (§ 4.1).
Propositional and modal logic reasoning in language models. Recent work has explored training and finetuning language models specifically for logical reasoning (Clark et al., 2021; Hahn et al., 2021; Tafjord et al., 2022). Our work differs in two key aspects: we evaluate general-purpose language models through prompting, a cost-efficient setup that has been widely adopted in recent years, and we focus on propositional and alethic modal logic rather than temporal (Hahn et al., 2021) or epistemic (Sileo and Lernould, 2023) logic.111We acknowledge that any logic that involves non-truth-functional operators, including but not limited to first-order logic, temporal logic, and epistemic logic, can be viewed as a modal logic; however, we adopt the most restrictive sense of modal logic (Ballarin, 2023) and use it interchangeably with alethic modal logic. Additionally, unlike studies comparing LLM and human performance on categorical syllogisms (Eisape et al., 2024, inter alia),222 We refer readers to Zong and Lin (2024) for a more comprehensive review of categorical syllogisms. we focus on hypothetical and disjunctive syllogisms with considerations of modality.
Human logic reasoning.
Research on human reasoning capabilities has informed studies of LLM logical reasoning. Eisape et al. (2024) compared LLM syllogistic reasoning with human behavior results (Ragni et al., 2019) under the framework of the Mental Models Theory (Johnson-Laird, 1983). Lampinen et al. (2024) found similar content effects in human and LLM reasoning, supporting the need to control for common-sense knowledge in benchmarks (§ 3.2). Belem et al. (2024) studied human and LLM perception of uncertainty at a lexical level. Compared to them, we focus on the modal logic reasoning process and contribute new behavioral data.
@@ 3 Dataset
We curate a dataset of natural-language multi-choice questions to measure the logical inference performance of LLMs (Figure 2). First, we use both propositional and modal logical forms as templates (§ 3.1). Then, we assign meanings (e.g., real-world interpretations or phrases with nonsensical words) to each propositional variable in the templates, and translate templates into natural-language Yes/No questions (§ 3.2).
@@ 3.1 Background: Propositional and Modal Logic
In propositional logic, a proposition can be evaluated as either true or false; however, this system can be overly simplistic when dealing with the complexity of real-world events. Consider the statement Alice is not eating., while it is true in a world where Alice is not eating, it may become false in a hypothetical possible world where Alice is indeed eating. This idea, known as possible world semantics (Kripke, 1959), provides a framework for expressing more nuanced statements about event possibilities, such as Alice may be eating and Alice must be eating. The former statement can be understood as there exists a possible world where Alice is eating, and the latter can be understood as in all possible worlds, Alice is eating.333 The possible world semantics, therefore, connects the notion of necessity and possibility to the universal and existential quantification () under first-order logic.
Normal modal logic (Kripke, 1963) follows this idea and extends propositional logic to reason about the necessity and possibility of events. A normal modal logic system can be formally described in the Backus–Naur form as
@@
Deduction and sequent.
Given a formula set as premises, if a deduction to a conclusion exists using axiom schemata and inference rules under the normal modal logic, we say the premises infer the conclusion, and the deduction can be represented as a logic sequent . If a formula set do not infer the conclusion, we denote it as and call it a non-entailment.
@@ 3.2 Translating Logic to Natural Language
An interpretation is a mapping from propositional variables to concrete meanings. For example, under the interpretation that is “Jane is eating apples” and is “John is eating oranges”, the logical formula becomes “Jane is eating apples or John is eating oranges.”
@@ 3.3 Involved Logical Forms
With the same interpretation, translates to
| Consider the following statements: |
| If Jane isn’t watching a show, then John is reading a book. |
| Jane isn’t watching a show. |
@@ 4.2.1 Analysis on Logic Sequents
Fixed effects. In line with Gonen et al. (2023) and McCoy et al. (2024), we find a negative correlation between perplexity and soft accuracy (); however, the correlation between them is weak (), which suggests the necessity of the complementary factors below in predicting LLM performance.
For different modalities and argument forms, we estimate their marginal means on soft accuracy (Figure 3), and perform pairwise hypothesis testing on the estimated coefficients (Table 2). The results generally align with the general observations on the full dataset. The only exception is that modus ponens (), instead of disjunctive syllogism (), appears to be the easiest argument form (i.e., the one with the highest soft accuracy) among all.
@@ 4.2.3 The Affirmation Bias over Modalities
One key argument of Dentella et al. (2023) is that large language models exhibit a bias towards affirming the claim, i.e., answering Yes more frequently than No.
We investigate this phenomenon by fitting a mixed-effects model
@@ 5 Human Experiments
The average human accuracy on each group is shown in the last row of Table 1.999Human responses are binary classes, so correct and incorrect responses are coded as and , respectively. Aligned with our LLM results (§ 4), on modalities the overall human results also show an accuracy order of (), and on argument forms, modus ponens () is the most accurately answered pattern.
@@
where is the binary accuracy of human responses, and Rt is the response time. The generalized mixed-effects model yields a marginal of yet a conditional , indicating a diverse response pattern across participants. The likelihood ratio test on the full model against the null model shows that only the effect of argument form is significant (, ). However, in accordance with the overall performance, we find modus ponens () has a significantly higher effect than other two valid argument forms. This confirms that logical forms can also have a significant impact on human reasoning accuracy, which is consistent with the LLM results, although the effect sizes are not the same.
@@ 6 Conclusion and Discussion
Probability and, relatedly, perplexity are often used as intrinsic evaluation metrics for language models. While Gonen et al. (2023) and McCoy et al. (2024) show that probability and perplexity, to some extent, correlates with LLM performance, literature in program synthesis with LLMs shows little correlation between probability and execution-based evaluation results (Li et al., 2022; Shi et al., 2022). This work does not necessarily contradict either line, but rather provides complementary factors for analyzing LLM performance.
@@
We argue that probability may have become an overloaded term in analyzing LLMs. Low probability may be due to one or more of the following non-exhaustive reasons: (1) the sentence is out of context, (2) the sentence is in ungrammatical language, or (3) the sentence is grammatical but has semantically awkward content (e.g., the mirror dataset in § 4.2.2), (4) the sentence contains rare content while being reasonable. We hypothesize that the probability of language models may not be essentially able to capture all these nuanced differences, and call for encoding and decoding algorithms—such as Meister et al. (2023)—that can better decompose the probability into finer-grained components.
Comparing humans and LLMs.
@@
What is our goal for building LLMs? Is it to achieve better performance on practical tasks or to build a more human-like model? Our results, together with Eisape et al. (2024), suggest that these two goals may not be perfectly aligned by revealing a mixture of similarity and discrepancy between LLMs and humans—for example, while LLMs exhibit higher benchmark performance than humans on our dataset and show the same argument form preferences with humans (Figures 3 and 7), they also show systematic biases that humans do not have (e.g., disfavoring the necessity modality, § 4.2.3). While there has been positive evidence of using LLMs as human models in psycholinguistic studies (Misra and Kim, 2024, inter alia), our results reaffirm that we should execute such approaches cautiously.
On the relation between modality and performance.
@@
Our results show that there is a significant difference in performance between necessity and possibility modalities, with the former much lower than the latter (Table 1). Part of the reason for this is that LLMs have a significant tendency to say “No” to necessity modality (Figure 6).
@@
On the one hand, our results extend the conclusion of Dentella et al. (2023), who found that LLMs generally respond positively—LLM behaviors may be significantly affected by finer-grained factors, including but not necessarily limited to the logical modality of the input. On the other hand, while LLMs systematically tend to answer “No” to questions in necessity modality, we do not find related evidence in human experiments, which leads to us hypothesize that such rejection bias comes from either the model architecture or the training strategies, such as the reinforcement learning with human feedback (RLHF; Ouyang et al., 2022) protocol. We leave this as an open question for future research.
Modal logic and theory of mind.
@@
Modality, in principle, encodes mental states and beliefs. The reasoning of beliefs also resonates with the theory of mind (Premack and Woodruff, 1978; Baron-Cohen et al., 1985, inter alia) and machine theory of mind (Rabinowitz et al., 2018; Ma et al., 2023, inter alia). Following the effort by Sileo and Lernould (2023) that uses epistemic modal logic to model the machine theory of mind, our work assesses the behaviors of LLMs on alethic modal logic, distantly revealing the future potential of LLMs in achieving the theory of mind.
@@
While we have verified that our data has a low perplexity ( under mistra-7b; much lower than that of the data by Wan et al. (2024), ), and, therefore, are similar enough to natural language utterances, the synthetic language cannot fully substitute natural language in daily life. Our dataset and analysis are not comprehensive enough to cover many nuanced examples that may appear in real communication, especially when context-dependent understanding is crucial to conveying communication goals.
@@
We do not foresee risk beyond the minimal risk posed by LLM evaluation work. We acknowledge that using LLMs in real-world scenarios could significantly impact human behaviors, raising the need for model transparency, safety, security, and interpretability. This work contributes to the understanding of We will open-source the synthetic logical reasoning dataset upon publication.
@@ A.2 Human Experiment Details
Participant consent.
We use the following language to obtain consent from participants, where our institution name is replaced with the Anonymous Institution to protect the anonymity of submission.
This study is part of a scientific research project at the Anonymous Institution. Your decision to complete this study is voluntary. There is no way for us to identify you. The only information we will have, in addition to your responses, is the demographic information you provided to Prolific and the time at which you completed the survey. The results of the research may be presented at scientific meetings or published in scientific journals. Clicking on the button below indicates that you are at least 18 years of age and agree to complete this study voluntarily. Press the button below to start the experiment.
@@
To mitigate the possible bias introduced by the dominant hand, we have the other half of the participants see instruction with reversed keys: In this study, you will be presented with two statements followed by a question. Your task is to answer either Yes or No to the question, based on the information provided in the statements. Please respond quickly and accurately by pressing "F" for No, and "J" for Yes.
@@ B.1 Considerations in Translating Logical Form to Natural Language
During the interpretation process, another key point is to assign independent interpretations to variables. Deciding the dependency also involves common sense knowledge. For example, consider the premises and . If we interpret and to proposition variables and , the two variables are possibly not independent. According to common sense, “Jane is not inside the house” () correlates with or is even equivalent to “Jane is out” (). Logically, ; however, with the extra premise given by common sense, people may conclude that .101010 This confounding factor affects the examples in Figure 10 of Holliday and Mandelkern (2024).

### v2

@@ Abstract
With the increasing interest in using large language models (LLMs) for planning in natural language, understanding their behaviors becomes an important research question. This work conducts a systematic investigation of LLMs’ ability to perform logical reasoning in natural language. We introduce a controlled dataset of hypothetical and disjunctive syllogisms in propositional and modal logic and use it as the testbed for understanding LLM performance. Our results lead to novel insights in predicting LLM behaviors: in addition to the probability of input (Gonen et al., 2023; McCoy et al., 2024), logical forms should be considered as important factors. In addition, we show similarities and discrepancies between the logical reasoning performances of humans and LLMs by collecting and comparing behavioral data from both.
@@ 1 Introduction
In this work, we investigate the logical reasoning capabilities of LLMs by assessing their performance on different logical forms. We curate a dataset of natural language statements and questions based on several logical forms in both propositional and modal logic, which is designed to mirror reasoning in daily communication. An example is shown in § 3.3. We then conduct a series of controlled experiments to analyze the performance of a set of LLMs on the dataset. Although our findings generally align with those by Gonen et al. (2023) and McCoy et al. (2024), who suggest that LLMs excel on examples with high probability, our results indicate that logical form, including but not limited to modalities and argument forms, is a crucial complementary factor in predicting the performance of LLMs (Figure 1). Additionally, with meaningful real-world interpretations, we find that:
@@
This paper is structured as follows. After reviewing related work (§ 2), we describe the dataset synthesis process (§ 3). We report the LLM reasoning results on our data (§ 4) and compare them with human performance (§ 5). We conclude by discussing the implications of our results and the limitations(§ 6).
@@ 2 Related Work
Logical reasoning benchmarks. Existing LLM logical reasoning benchmarks (Liu et al., 2020; Han et al., 2024, inter alia) focus on complex, multi-hop reasoning problems with manually annotated problems, making cross-problem comparisons challenging. Recent work has introduced benchmarks with synthesized natural-language questions using predefined logical formulas and substitution rules (Saparov and He, 2022; Saparov et al., 2023; Parmar et al., 2024; Wan et al., 2024, inter alia). Compared to them, our work uniquely incorporates modal logic, which has been largely unexplored in existing benchmarks—while Holliday et al. (2024) present a case study, our approach offers two key advances: controlled knowledge bias in logic interpretations (§ 3.3) and a more rigorous statistical evaluation framework (§ 4.1).
Propositional and modal logic reasoning in language models. Recent work has explored training and finetuning language models specifically for logical reasoning (Clark et al., 2021; Hahn et al., 2021; Tafjord et al., 2022). Our work differs in two key aspects: (1) we evaluate general-purpose language models through prompting, a cost-efficient setup that has been widely adopted in recent years, and we focus on propositional and alethic modal logic rather than temporal (Hahn et al., 2021) or epistemic (Sileo and Lernould, 2023) logic; 111Technically, any logic that involves non-truth-functional operators, including first-order logic, temporal logic, and epistemic logic, can be viewed as a modal logic; however, we adopt the most restrictive sense of modal logic (Ballarin, 2023) and use it interchangeably with alethic modal logic. (2) unlike studies comparing LLM and human performance on categorical syllogisms (Eisape et al., 2024, inter alia),222 We refer readers to Zong and Lin (2024) for a more comprehensive review of categorical syllogisms. we focus on hypothetical and disjunctive syllogisms with considerations of modality.
Human logic reasoning. Work on human reasoning capabilities has informed studies of LLM logical reasoning: Eisape et al. (2024) compared LLM syllogistic reasoning with human behavior results (Ragni et al., 2019) under the framework of the Mental Models Theory (Johnson-Laird, 1983); Lampinen et al. (2024) found similar content effects in human and LLM reasoning, supporting the need to control for common-sense knowledge in benchmarks (§ 3.2); Belem et al. (2024) studied human and LLM perception of uncertainty at a lexical level. Compared to them, we focus on the propositional and modal logic reasoning process and contribute new behavioral data.
@@ 3 Dataset
We curate a dataset of natural-language multi-choice questions to measure the logical inference performance of LLMs. Starting from propositional and modal logical forms as templates (§ 3.1), we assign meanings (e.g., real-world interpretations) to each variable and translate templates into natural-language Yes/No questions (§ 3.2). A subsidiary visualization of the process is shown in Figure 2.
@@ 3.1 Background: Propositional and Modal Logic
In propositional logic, a proposition can be evaluated as either true or false; however, this system can be overly simplistic when dealing with the complexity of real-world events. Consider the statement Alice is not eating, while it is true in a world where Alice is not eating, it may become false in a hypothetical possible world where Alice is indeed eating. This idea, known as possible world semantics (Kripke, 1959), provides a framework for more nuanced statements about event possibilities, such as Alice may be eating and Alice must be eating. The former statement can be understood as there exists a possible world where Alice is eating, and the latter can be understood as in all possible worlds, Alice is eating.333 The possible world semantics, therefore, connects the notion of necessity and possibility to the universal and existential quantification () under first-order logic. Normal modal logic (Kripke, 1963) formalizes this idea and extends propositional logic to reason about event necessity and possibility. In the Backus–Naur form, a normal modal logic system can be written as
@@
Deduction and sequent. Given a formula set as premises, if a deduction to a conclusion exists using axiom schemata and inference rules under the normal modal logic, we say the premises infer the conclusion, and the deduction can be represented as a logic sequent . If a formula set do not infer the conclusion, we denote it as and call it a non-entailment.
@@ 3.2 Translating Logic to Natural Language
An interpretation maps propositional variables to concrete meanings. For example, under the interpretation that is “Jane is eating apples” and is “John is eating oranges”, the logical formula becomes “Jane is eating apples or John is eating oranges.”
@@ 3.3 Involved Logical Forms
With the same interpretation, ’s translation of the first statement is If Jane isn’t watching a show, then John is reading a book.
@@ 4.2.1 Analysis on Logic Sequents
| Hypothesis | -value |
Fixed effects. In line with Gonen et al. (2023) and McCoy et al. (2024), we find a negative correlation between perplexity and soft accuracy (); however, the correlation is weak (), which suggests the necessity of the complementary factors below in predicting LLM performance.
For different modalities and argument forms, we estimate their marginal means on soft accuracy (Figure 3) and perform pairwise hypothesis testing on the estimated coefficients (Table 2). The results generally align with the general observations on the full dataset. The only exception is that modus ponens (), instead of disjunctive syllogisms (), appears to be the easiest argument form (i.e., the one with the highest soft accuracy) among all.
@@ 4.2.3 The Affirmation Bias over Modalities
One key argument of Dentella et al. (2023) is that large language models exhibit a bias towards affirming the claim, i.e., answering Yes more frequently than No. We investigate this phenomenon by fitting a mixed-effects model
@@ 5 Human Experiments
The average human accuracy on each group is shown in the last row of Table 1.999Human responses are binary classes, so correct and incorrect responses are coded as and , respectively. Aligned with our LLM results (§ 4), on modalities, the overall human results also show an accuracy order of (), and on argument forms, modus ponens () is the most accurately answered pattern.
@@
where is the binary accuracy of human responses, and Rt is the response time. The generalized mixed-effects model yields a marginal of yet a conditional , indicating a diverse response pattern across participants. The likelihood ratio test on the full model against the null model shows that only the effect of argument form is significant (, ). However, in accordance with the overall performance, we find modus ponens () has a significantly higher effect than the other two valid argument forms. This confirms that logical forms can also have a significant impact on human reasoning accuracy, which is consistent with the LLM results, although the effect sizes are not the same.
@@ 6 Conclusion and Discussion
Probability in language models. Probability and perplexity are often used as intrinsic evaluation metrics for language models. While Gonen et al. (2023) and McCoy et al. (2024) show that probability and perplexity correlate well with LLM performance, literature in program synthesis with LLMs shows little correlation between probability and execution-based evaluation results (Li et al., 2022; Shi et al., 2022). This work does not necessarily contradict either line but rather provides complementary factors for analyzing LLM performance.
@@
We argue that probability may have become an overloaded term in analyzing LLMs. Low probability may be due to one or more of the following non-exhaustive reasons: (1) out-of-context content, (2) ungrammatical language, or (3) grammatical but semantically awkward content (cf. the mirror dataset in § 4.2.2), (4) reasonable but rare content. We hypothesize that the probability of language models may not be essentially able to capture all these nuanced differences, and call for encoding and decoding algorithms—such as Meister et al. (2023)—that can better decompose the probability into finer-grained and explainable components.
@@
Comparing humans and LLMs. What is our goal for building LLMs? To achieve better performance on practical tasks or to build a more human-like model? Our results, together with Eisape et al. (2024), suggest that these two goals may not be perfectly aligned by revealing a mixture of similarity and discrepancy between LLMs and humans—for example, while LLMs exhibit higher benchmark performance than humans on our dataset and show the same argument form preferences with humans (Figures 3 and 7), they also show systematic biases that we do not find significant in human reasoning (e.g., disfavoring the necessity modality, § 4.2.3). While there has been positive evidence of using LLMs as human models in psycholinguistic studies (Misra and Kim, 2024, inter alia), our results suggest executing such approaches cautiously.
@@
On the relation between modality and performance. Our results show that there is a significant difference in performance between necessity and possibility modalities, with the former much lower than the latter (Table 1). Part of the reason for this is that LLMs have a significant tendency to say “No” to the necessity modality (Figure 6).
@@
On the one hand, our results extend the conclusion of Dentella et al. (2023) that LLMs generally respond positively—LLM behaviors may be significantly affected by finer-grained factors, including but not necessarily limited to the modality involved in the input. On the other hand, while LLMs systematically tend to answer “No” to questions in necessity modality, we do not find related evidence in human experiments, which leads us to hypothesize that such rejection bias comes from either the model architecture or the training strategies, such as the reinforcement learning with human feedback (RLHF; Ouyang et al., 2022) protocol. We leave this as an open question for future research.
@@
Modal logic and theory of mind. Modality, in principle, encodes mental states and beliefs. The reasoning of beliefs also resonates with the theory of mind (Premack and Woodruff, 1978; Baron-Cohen et al., 1985, inter alia) and machine theory of mind (Rabinowitz et al., 2018; Ma et al., 2023, inter alia). Following the effort by Sileo and Lernould (2023) that uses epistemic modal logic to model the machine theory of mind, our work assesses the behaviors of LLMs on alethic modal logic, distantly revealing the future potential of LLMs in achieving the theory of mind.
@@
While we have verified that our data has a low perplexity ( under mistral-7b; much lower than that of the data by Wan et al. (2024), ), and, therefore, are similar enough to natural language utterances, the synthetic language cannot fully substitute natural language in daily life. Our dataset and analysis are not comprehensive enough to cover many nuanced examples that may appear in real communication, especially when context-dependent understanding is crucial to conveying communication goals.
@@
This work contributes to the understanding of LLMs. We do not foresee risk beyond the minimal risk posed by LLM evaluation work. We acknowledge that using LLMs in real-world scenarios could significantly impact human behaviors, raising the need for model transparency, safety, security, and interpretability. We will open-source the synthetic logical reasoning dataset upon publication.
@@ A.2 Human Experiment Details
To mitigate the possible bias introduced by the dominant hand, we have the other half of the participants see instruction with reversed keys:
In this study, you will be presented with two statements followed by a question. Your task is to answer either Yes or No to the question, based on the information provided in the statements. Please respond quickly and accurately by pressing "F" for No, and "J" for Yes.
@@ B.1 Considerations in Translating Logical Form to Natural Language
During the interpretation process, another key point is to assign independent interpretations to variables. Deciding the dependency also involves common sense knowledge. For example, consider the premises and . If we interpret and to proposition variables and , the two variables are possibly not independent. According to common sense, “Jane is not inside the house” () correlates with or is even equivalent to “Jane is out” (). Logically, ; however, with the extra premise given by common sense, people may conclude that .101010 This confounding factor affects the examples in Appendix C.1.12 of Holliday et al. (2024).